OpenStudy (goformit100):

I earlierly have posted this question, but bymistake I closed it... So please let me allow to get the answer again The depth d(x), in feet, of a circular swimming pool at a point x feet from the edge is represented by the function below. The function d of x equals 1 plus one-eighth x] If the depth varies from 1 foot to 4 feet, which inequality describes the domain of the function? 0 ≤ x ≤ 8 0 ≤ x ≤ 24 0 ≤ x ≤ 40 1 ≤ x ≤ 8

OpenStudy (mathstudent55):

\( d(x) = 1 + \dfrac{1}{8} x \) When \(d(x) = 1\), \( 1 = 1 + \dfrac{1}{8} x \) \(0 = \dfrac{1}{8} x \) \(0 = x\) When \(d(x) = 4\), \( 4 = 1 + \dfrac{1}{8} x \) \(3 = \dfrac{1}{8} x \) \(24 = x\) For d(x) to be between 1 and 4, x is between 0 and 24, and that is domain.

OpenStudy (goformit100):

What's the correct option then ?

OpenStudy (goformit100):

@mathstudent55

OpenStudy (mathstudent55):

I wrote above: "x is between 0 and 24, and that is domain." Don't you see an option with 0 and 24 in it?